OpenStudy (anonymous):

hey guys i have a math final tomorrow and i do not know how to solve like this problem : evaluate the definite integral : integral ( 1 to 2 ) 24(x^2-6)^5x dx = .. please could someone tell me how to start solving these problems i have tried many times but i keep getting wrong answer :(

OpenStudy (callisto):

Is the question\[\int_1^224(x^2-6)^5xdx\]

OpenStudy (anonymous):

yes , but how can i solve it

OpenStudy (callisto):

Let u = x^2 -6 du = ... ?

OpenStudy (anonymous):

so du = 2x ?

OpenStudy (anonymous):

then 1/2x du = dx ?

OpenStudy (callisto):

1/2x du = dx <- yes

OpenStudy (callisto):

Next step is to change the limit. Remember we let u = x^2 -6 When x = 1, u = 1^2-6 = -5 When x = 2, u = ...?

OpenStudy (anonymous):

-2

OpenStudy (callisto):

\[\int_1^224(x^2-6)^5xdx\]\[=\frac{1}{2}\int_1^224(x^2-6)^5(2xdx)\]Note: 2xdx = du. Now, we do the u-sub: \[=\frac{1}{2}\int_{-5}^{-2}24(u)^5(du)\]So far so good?

OpenStudy (anonymous):

yess

OpenStudy (callisto):

Now, do the integration :P

OpenStudy (anonymous):

thats it ?

OpenStudy (anonymous):

omg thank you so much ^^

OpenStudy (callisto):

You're welcome! Remember to do the integration LOL.

OpenStudy (anonymous):

Yup I will .. :P